OpenStudy (anonymous):

Solve the problem as an inequality. A parallelogram has a base of 15 and an area which is less than or equal to 45. What is its heigh Solve the problem as an inequality. A parallelogram has a base of 15 and an area which is less than or equal to 45. What is its heigh @Mathematics

OpenStudy (anonymous):

height

OpenStudy (hoblos):

\[h \le 3\]

OpenStudy (anonymous):

X < 3 A = BH = 15 * H A <= 45 15H <= 45 H <= 3